OpenStudy (anonymous):

the sum of two consecutive integers is -43. Find the integers

OpenStudy (eyust707):

...

OpenStudy (anonymous):

?

OpenStudy (eyust707):

find the two integers?

OpenStudy (anonymous):

yes, sorry

OpenStudy (eyust707):

an interger is any number like ...-12,-11,-10... 0,1,2,3,4...

OpenStudy (anonymous):

I got 21 and 22 but i dont know how it would be negative

OpenStudy (shadowlegendx):

Do you have any answer choices?

OpenStudy (eyust707):

they have to be "consecutive" which means they are right next to each other on the number line... like 3 and 4. or -21 and -20

OpenStudy (anonymous):

No its fill in the two blanks:(

OpenStudy (shadowlegendx):

I got: -21 - 22 = -43

OpenStudy (eyust707):

-21 and -22 would work -21 + (-22) = -43

OpenStudy (anonymous):

Ohhhhhh Lol Thanks Soo Much :)

OpenStudy (eyust707):

they are basically testing your vocab here more than your math ;)

OpenStudy (shadowlegendx):

Yeah, if you simplify you get: -21 + (-22) = -43

OpenStudy (anonymous):

what about "a novel is discounted 30%. The sale price is $20.65 what was the original price? "